# Space Group Informed Transformer for Crystalline Materials Generation

### v1

@@ III Results and Evaluations
We train the CrystalFormer using the MP-20 dataset [11] which is a popular dataset that represents a majority of experimentally known crystalline materials at ambient conditions with no more than 20 atoms in the primitive unit cell. The training dataset contains 27136 crystal structures. The subdivision of the training samples according to the space group has greatly reduced the number of samples in each space group category. On top of that, the distribution of training samples is quite uneven among the space groups, which reflects the imbalance distribution of crystals over space groups in Nature [30]. In fact, there is no training data in 61 out of 230 space groups as shown in Fig. 6. Nevertheless, we still employ the MP-20 as the training set so that the performance of the model can be more easily gauged with the others in the literature.
Figure 2 shows a breakup of the learning curves for the Wyckoff position, chemical species, fractional coordinates, and lattice parameters. Next, we will select the model checkpoint with the lowest total validation loss and examine its learned features. Then, we will use the model to generate 1000 crystal samples for each of the 230 space groups and evaluate their validity, stability, and novelty. These results highlight the benefits of built-in space group symmetries in the model and point to possible future applications of the CrystalFormer model.
III.1 Learned features
Figure 3 visualizes the cosine similarity of the learned vector embedding of the chemical species. Red colors in the figure indicate similar chemical species identified by the model. One clearly sees the periodicity shows up as off-diagonal stripes. Moreover, there are visible clusters for Lanthanide elements (La-Lu). The plot also suggests the similarity between the lanthanides and other rare-earth elements (Y and Sc). Being able to learn chemical similarities from data [42, 43, 44, 45, 46, 19] is an encouraging signal that the model is picking up atomic physics to be able to generate reasonable crystal structures. In particular, the features shown in Fig. 3 is strikingly similar to the similarity map constructed purposely based on substitution pattern [42] which was later used for substitution-based material discovery [47].
Figure 4 presents the histogram of the total number of atoms in the conventional unit cell for several space groups. One sees a nice agreement between the atom number distribution in the test dataset and the generated samples. In addition, it appears that space group is the key latent variable that decomposes the multi-modal atom number distribution of crystals. This is understandable because the number of atoms is determined by the sum of the multiplicities of occupied Wyckoff positions. Therefore, the space group symmetry is a key control variable for the atom number distribution. Building Wyckoff positions information into the CrystalFormer model architecture removes the necessities of querying the training data to find out the number of atoms for a targeted space group [16].
Figure 5 shows heat maps of Wyckoff positions and chemical species for the space group (No. 225). First of all, one sees that most atoms occupy those special Wyckoff positions with high symmetries. Moreover, the heat map reveals interesting solid-state chemistry knowledge as it suggests where each atom tends to appear in a crystal. For example, there are vertical stripes at the locations of inert elements (He, Ne, Ar…) as they are rare in crystalline materials. Moreover, one sees that oxygen and halogen elements (F, Cl, Br, I) appear quite often in the Wyckoff position ”24e”, which is a consequence of their electronegativities. Overall, we see the CrystalFormer has learned these key motifs for generating crystalline materials. On the other hand, one sees that several Wyckoff locations of the hydrogen are missing in the generated samples compared to the test dataset. We believe that is due to that the hydrogen element takes only about in the training data for the space group. Collecting more data with better coverage of elements will be crucial to further boost the performance of the current model.
There has been a long history of summarizing empirical or statistical chemistry rules encoded in materials data and then use them to instruct search of crystal structures [48, 49, 50, 51, 42, 52]. Our analysis shows that CrystalFormer also ingests a number of chemical intuitions in the training data for generative modelling. Since training procedure compresses the chemistry knowledge into a neural network model with tractable likelihoods, it is possible to employ CrystalFormer in the Monte Carlo or evolution strategy search of crystal structures, besides as a probabilistic generative model.
III.2 Validity and novelty of generated samples
Figure 6 illustrates the structure and compositional validity of generated samples across all 230 space groups. Following the Ref. [53], a structure meets the validity criteria if the shortest atomic distance exceeds , a lenient standard. Composition validity requires charge neutrality as computed by SMACT [54]. This is, however, an overly stringent criterion since the composition validity of the training set is only around by this measure [11]. Note that the CrystalFormer is able to generate reasonable samples even for those space groups without any training data.
| Space group | Crystal system | Training samples | Validity () | |
| Struc. | Comp. | |||
| 2 | Triclinic | 676 | 83.10 | 83.0 |
| 12 | Monoclinic | 1273 | 87.70 | 81.80 |
| 62 | Orthorhombic | 1187 | 95.50 | 87.20 |
| 139 | Tetragonal | 1233 | 97.70 | 83.40 |
| 166 | Trigonal | 1076 | 98.50 | 85.0 |
| 194 | Hexagonal | 1129 | 99.40 | 89.90 |
| 225 | Cubic | 3960 | 99.60 | 93.50 |
| 1 | Triclinic | 27136 | 91.40 | 80.20 |
| Autoregressive models | ||||
| PGSchNet [55] | 99.65 | 75.96 | ||
| LM-CH (character-level tokenization) [18] | 84.81 | 83.55 | ||
| LM-AC (atom coordinate-level tokenization) [18] | 95.81 | 88.87 | ||
| Crystal-LLM [20] | 96.5 | 86.3 | ||
| Diffusion models | ||||
| CDVAE [11] | 100.0 | 86.70 | ||
| DiffCSP [13] | 100.0 | 83.25 | ||
| DiffCSP++ [26] | 99.94 | 85.12 | ||
| UniMat-Large [15] | 97.2 | 89.4 | ||
Table 2 reports the validity of generated samples for selected space groups in each of the seven crystal systems. To ensure the numbers are representative, we chose the space group to be the one with the most training data for each crystal system. One sees that the model performs better for more symmetric space groups. This is a nice feature that complements existing crystal generative models, which mostly have difficulties in generating highly symmetric structures. As a reference, we also list the validity of the generated samples suppose one treats the crystals as if they are all in the space group (No. 1) with only translational symmetry. One sees the structure validity scores of space group improves compared to the one shown in Figure 6 due to increased training samples.
The second part of Table 2 shows reference results in the literature for the same validity test. In principle, the performance of the present model should fall back to the language model approaches [18, 20]. The remaining gap may be due to details such as the including the header line in the crystallographic information files (CIF), specific sampling strategy of language models, or the additional post-selection of samples [20]. In the table, the DiffCSP++ [26] is the only alternative model that exploits the space group symmetry in the generation process. However, different from us, the DiffCSP++ model does not predict Wycokff position as a part of generation process. Therefore, DiffCSP++ needs to search for template structure in the training set for generation, which may limit its generality. Besides works listed in Table 2 that reported validity scores in a comparable settings, Ref. [19] has conditioned the generation of CIF on the space group symbols in a language model setting. Ref. [16] considered space group conditioned crystal generation using a fine-tuned generative model with space group labels. Neither approach provides exact constraints on the space group, which could yield problematic structures for large systems and highly symmetric space groups. Ref. [23] considered generating symmetric crystals in their Wyckoff representation. However, the model does not consider fractional coordinates and lattice parameters, so it requires a subsequent computational search to completely determine the crystal structure. Sec. IV.2 contains a more in-depth discussion of the relationship of the present approach with related works.
Figure 7(a) shows the validity of generated samples as a function of sampling temperature. One clearly sees that reduced temperature increases the validity of samples at the cost of reducing the diversity [20]. Figure 7(b) shows the novelty and uniqueness evaluated on generated samples with temperature. Novelty quantifies the proportion of new structures in the generated samples that were unseen in the training dataset. Uniqueness represents the percentage of distinct, non-redundant structures among the generated samples [56]. One sees that across different temperatures the uniqueness remains high, indicating the model does not collapse to a mode that produces duplicated samples. On the other hand, the model produces close to novel material as temperature increases, which nicely demonstrates modal covering behavior of the maximum likelihood estimation training [57]. Having a model distribution broader than the span of the dataset is crucial for material discovery.
III.3 Stability of generative samples
Figure 8(a) shows the histogram of the log-likelihood of generated and test samples show nice agreement. We also visualize structures of a few generated samples which deems to be very likely, typical, and unlikely according to their log-likelihoods. Interestingly, we did not observe a clear correlation between the model likelihood and energy of the structure.
Fig. 8(b) shows the energy above hull based on energies calculated with density functional theory (DFT). One sees the distributions of unrelaxed and relaxed structures are very similar, suggesting that the generated structures are already very close to local minima of DFT calculations. For the given computational budget detailed in the Appendix B, we manage to reach convergence for 973 out of 1000 structures in the DFT relaxation. All of the structures retains the same space group symmetry after the DFT relaxation 222Due to merging of Wyckoff positions that are occupied with the same elements, 26 structures actually belong to the space group (No. 221).. With its default parameters, StructureMatcher from pymatgen [60] is able to match all 973 structures before and after relaxation. The average root mean squared displacement (RMSD) normalized by computed for these matched structures is 0.0061, which confirms the generated samples are indeed very close to DFT local minima. Figure 8(b) shows a pronounced peak around eV/atom showing the model indeed generates a large fraction of candidates of metastable materials [58]. Finally, we found 21 relaxed samples are stable by the criteria and summarize them in Table S2 of appendix B. Most of the found stable materials only have atoms occupying the first few Wyckoff positions. Among these 21 structures, we found 10 are new crystalline materials that are not contained in the Materials Project database [61].
IV Discussions
IV.1 Applications
With the abilities demonstrated in Sec. III, we believe that CrystalFormer is ready to fit into the existing materials discovery workflow in the following two ways.
IV.1.1 Initialization for structure search
Crystal structure search relies heavily on utilizing space group symmetries. For example, it is a common practice for crystal structure prediction software [62, 63, 64, 65, 66] and crystal structure search [67, 68] to place atoms at Wyckoff positions to prepare symmetric initial states for subsequent optimization.
However, such initialization approach faces combinatorial difficulty as the number of chemical species and atoms in the unit cell grow. The CrystalFormer is ready to act as a drop-in replacement of random structure initialization. In this way, one bypasses the curse-of-dimensionality of exact enumeration with a data-driven probabilistic approach. Moreover, the ability of CrystalFormer to generate diverse and close to metastable structures can greatly reduce computational costs of downstream ab initio calculations.
IV.1.2 Mutation of existing structures
Mutatation of known crystals is a prominent approach to materials discovery. For example, one can employ machine machine-learned force field to relax crystal structures [69, 5, 70, 71] after element substitutions. In the lens of generative modeling, the machine learning force field can be regarded as the energy-based model or Boltzmann machines. A potential drawback of exploring materials space with an energy-based model is the slow mixing or even ergodicity issue posed by the rough landscape of the potential energy surface. In this sense, element substitutions provide a variety of initial seeds, compensating the limitation potential energy surface-based exploration.
@@ IV.1.2 Mutation of existing structures
Having an alternative measure of crystal likelihood other than the potential energy surface opens a way to employ the model likelihood as a guide for structure search. For example, with a trained CrystalFormer, one can perform Markov chain Monte Carlo (MCMC) random walk in the crystalline materials starting from an existing crystal structure. At each step of the random walk, one proposes an element substitution, atom shift, or lattice deformation and accept or reject the proposal according to the model likelihood. In the long time limit, one will sample from the model distribution. However, during the burn-in phase of such MCMC sampling, the generated samples will be similar to the starting material, which may be a desired feature in certain cases.
IV.2 Related works

### v2

@@ II.2 Training
In the present work, we train the CrystalFormer using the MP-20 dataset [11] which is a popular dataset that represents a majority of experimentally known crystalline materials at ambient conditions with no more than 20 atoms in the primitive unit cell. The training dataset contains 27136 crystal structures. The subdivision of the training samples according to the space group has greatly reduced the number of samples in each space group category. On top of that, the distribution of training samples is quite uneven among the space groups, which reflects the imbalanced distribution of crystals over space groups in nature [30]. In fact, there is no training data in 61 out of 230 space groups as shown in Fig. S2. Nevertheless, we still employ the MP-20 as the training set so that the performance of the model can be more easily gauged with the others in the literature, see appendix B. Figure 2 shows a breakup of the learning curves for the Wyckoff position, chemical species, fractional coordinates, and lattice parameters. We select the model checkpoint with the lowest total validation loss to generate crystal samples.
II.3 Sampling
To sample crystals from the CrystalFormer, one needs to specify a space group number and a list of possible chemical elements. The CrystalFormer samples the atoms one by one, starting from more symmetric specific positions with lower multiplicities till less symmetric general positions with larger multiplicities. We use the information of the space group and Wyckoff letter to control the sampling of fractional coordinates. By applying the symmetry projection to the sampled fractional coordinate, one rectifies it and ensures the generated fractional coordinates are compatible with the Wyckoff positions. One can also mask out the logits of chemical species so that only a number of selected elements will be sampled. The number of symmetrically inequivalent atoms may fluctuate in the sampling procedure. Once one has sampled a padding atom, the model predicts the lattice parameters under the space group constraint. Moreover, we introduce a temperature parameter in the sample distribution . With we will draw samples from a sharper distribution, while gives more diversity in the generated samples. In the present paper, we will generate crystals using temperature unless mentioned explicitly.
@@ II.3 Sampling
Besides autoregressive sampling, one can also perform Markov chain Monte Carlo (MCMC) sampling based on the likelihood Eq. (1) of the CrystalFormer. MCMC sampling of crystals is a way to walk through the crystalline materials starting from an existing crystal structure. At each step of the random walk, one proposes an element substitution, atom position shift, or lattice deformation to change the crystal from to , then accepts or rejects the proposal according to the model probability according to the Metropolis acceptance rate . MCMC sampling is particularly useful for incorporating additional constraints or guidance in the sampling procedure. Moreover, during the burn-in phase of such MCMC sampling, the generated samples will be similar to the starting material, which may be a desired feature in certain cases.
III CrystalFormer learns chemical intuition by compressing materials database
Nature tends to favor symmetrical crystal structures. Crystallographic space groups capture and quantify this inductive bias of nature, thereby significantly simplifying the spaces of crystal materials. In light of the space group symmetries, crystals also have an unexpected yet natural sequential and discrete representation, which derives from two tables in nature: the periodic table of elements determined by quantum mechanics and the table of Wyckoff positions of the 230 space groups determined by group theory. To construct a certain crystal, we only need to select atoms from the periodic table and place them sequentially into the Wyckoff positions in the unit cell. In this crystal language, the “word order” is determined by the alphabetical order of Wyckoff letters, the “grammar” corresponds to the solid-state chemistry rules, and the “synonyms” represent interchangeable elements (Sec. III.1), the “sentence length” correspond to atom number in the unit cell (Sec. III.2), and the “idioms” correspond to common chemical coordination (Sec. III.3).
CrystalFormer employs an autoregressive transformer to learn the crystal language, thereby exploring yet-to-be-discovered crystalline materials. It compresses and internalizes the crystal materials database, expressing solid-state chemical knowledge through neural network parameters; reflecting the associative ability of material space through neural network activations; and describing chemical intuition through the model probability (Sec. III.4). Similar to generative models used for generating text, images, and videos, CrystalFormer can directly generate “realistic” crystal materials. However, rather than worrying about the fake contents of AI-generated media, these AI-generated crystal materials could potentially be synthesized and be useful to human civilization.
Next, we will inspect the learned features and sample statistics of the model to build up an understanding of the CrystalFormer. We carry out inspectations for a few selected space groups. The findings are neverthelss general. These findings not only provide confidence a bit of understandings of the model, but also direct us to the suitable applications of CrystalFormer.
III.1 Atom embeddings and chemical similarity
Figure 3 visualizes the cosine similarity of the learned vector embedding of the chemical species. Red colors in the figure indicate similar chemical species identified by the model. One sees the chemical similarity within groups of elements show up as off-diagonal red stripes. Moreover, there are visible clusters for Lanthanide elements (La-Lu). The plot also suggests the similarity between the lanthanides and other rare-earth elements (Y and Sc). The features shown in Fig. 3 are strikingly similar to the similarity map constructed purposely based on substitution pattern [46, 47] which was later used for substitution-based material discoveries [48, 5]. In the context of language modeling, the chemical similarities correspond to synonyms of chemical species tokens. Having the ability to learn them from data [46, 47, 49, 50, 51, 52, 53, 19] is an encouraging signal that the model is picking up atomic physics to be able to generate reasonable crystal structures.
III.2 Atom number distributions
The number of atoms corresponds to the length of non-padding atoms in the sequential representation, which is captured well by CrystalFormer. Figure 4 presents the histogram of the total number of atoms in the conventional unit cell for several space groups. One sees a nice agreement between the atom number distribution in the test dataset and the generated samples. In addition, it appears that space group is the key latent variable that decomposes the multi-modal atom number distribution of crystals. This is understandable because the number of atoms is determined by the sum of the multiplicities of occupied Wyckoff positions. Therefore, the space group symmetry is a key control variable for the atom number distribution. Incorporating Wyckoff positions information into the CrystalFormer model architecture removes the necessity of querying the training data to find out the number of atoms for a targeted space group [16].
Recently, Ref. [54] reports an abundance of inorganic compounds whose primitive unit cell contains a number of atoms that is a multiple of four. There are different ways to reason about the observed ”rule of four” depending on one’s view of how a crystal is formed. For example, one can often break inorganic solids into polyhedra as building blocks. Otherwise, Ref. [55] considers the most probable values of the number of atoms in a formula unit and the number of formula units per primitive cell. In line with the discussion here, the ”rule of four” is the combination of three factors 1) the distribution of crystalline materials among space groups [30]; 2) the distribution of atoms in Wyckoff positions [33] of a given space group; and 3) the multiplicities of Wyckoff positions and multiplicities of conventional versus cells. The first two are statistical rules determined by the inter-atomic interactions while the third one is a mathematical fact of space group theory. In the end, the point we want to make is that the design of CrystalFormer and its associated crystal representation allow it to learn the ”rule of four” and many other to-be-discovered ”rules”, which manifest themselves as marginal statistics of learn probability distribution. Most importantly, CrystalFormer will utilize these ”empirical rules” when generating novel yet reasonable crystal samples.
III.3 Wyckoff-Atom gram
Figure 5 shows heat maps of Wyckoff positions and chemical species for the space group (No. 225). The heat map is analogous to bigram frequency statistics in language modeling. In the present context, it reveals interesting solid-state chemistry knowledge related to where each atom tends to appear in a unit cell. First of all, one sees that most atoms occupy special Wyckoff positions (Wyckoff letters at the beginning of the alphabet) with higher site symmetries. The distribution of generated data is in agreement with test data and recent statistics [33]. Moreover, there are vertical blanks at the locations of inert elements (He, Ne, Ar…) as they are rare in crystalline materials. Lastly, one sees that oxygen and halogen elements (F, Cl, Br, I) appear quite often in the Wyckoff position ”24e”, which is a consequence of their electronegativities [30]. Overall, we see the CrystalFormer has learned these key motifs for generating crystalline materials. Nevertheless, one sees that several Wyckoff locations of the hydrogen are missing in the generated samples compared to the test dataset. We believe that is due to that the hydrogen element takes only about in the training data for the space group. Collecting more data with better coverage of elements will be crucial to further boost the performance of the current model.
Along the same line of thoughts, coordination polyhedra [56] and lattice structure [57] manifest themselves as higher-order n-gram correlations of Wyckoff position and atom species in the crystal language, which will be captured by the CrystalFormer. There has been a long history of mining empirical chemistry rules encoded in materials data and then using them to instruct the search of crystal structures [58, 59, 60, 61, 47, 62]. Our analysis shows that CrystalFormer ingests chemical intuition, be it speakable or unspeakable, in the training data for generating new materials.
III.4 Crystal likelihoods
CrystalFormer compresses chemistry knowledge stored in the material dataset into its parameters. In addition to generating crystal samples, CrystalFormer can also compute the likelihoods of crystals via Eq. (1). Therefore, it is possible to employ CrystalFormer in the Monte Carlo search of crystal structures besides sampling materials directly as a generative model.
Figure 6 shows the agreement of the likelihoods of generated samples and samples in the test dataset. We also visualize structures of a few generated samples which are deemed to be very likely, typical, and unlikely according to their likelihood values. We have checked that likelihood is related to the energy of the crystal by locally perturbing the fractional coordinates and lattice parameters. However, we did not observe a correlation between the likelihood of these crystals and their energies on a global scale. We envision the landscape of likelihoods is much less rough compared to the potential energy surface of crystalline materials. Intuitively, it means that the Crystalformer compresses the materials space into a more compact space without many holes that correspond to infeasible high energy states. Therefore, likelihood-based exploration of the crystal space discussed in Sec. II.3 can be more efficient compared to traditional sampling approaches based on the physical energy functions.
IV Applications
We now move on to the practical applications of CrystalFormer to materials discovery and design. Compared to many existing materials generation models, CrystalFormer offers precise control over space group symmetry and enables efficient computation of model likelihood. These unique features open a wide range of possibilities for integrating it with existing computational software and machine-learning models in a flexible way as we demonstrate below. For these applications, we have excluded radioactive elements from the samples [32].
IV.1 Symmetry-conditioned random structure initialization
Crystal structure prediction has long been the dream of solid-state chemistry and computational material science researchers [63]. Typical crystal structure prediction workflow consists of two steps. First, one randomly initializes a batch of diverse crystal structures as candidates. Second, one optimizes the crystal structures via local and global optimization strategies. Utilizing space group symmetries plays a crucial role in both steps, as symmetry enlarges the span of the energy distribution [64, 65, 66] and reduces the search space.
It is a common practice for crystal structure prediction software [67, 68, 69, 65, 70, 66] and structure search [71, 72, 73] to exploit space group symmetry in the crystal structure initialization. However, such an initialization approach faces combinatorial difficulty as the number of chemical species and atoms in the unit cell grows. The CrystalFormer is ready to act as a drop-in replacement of random structure initialization for crystal structure prediction. In this way, one bypasses the curse-of-dimensionality of exact enumeration [65] with a data-driven probabilistic approach. Moreover, the ability of CrystalFormer to generate diverse and near-stable structures can greatly reduce the computational costs of downstream optimizations.
We select seven space groups (No. 2), (No. 12), (No. 62), (No. 139), (No. 166), (No. 194), and (No. 225) as representatives of the seven crystal systems. We randomly generate 100 crystals for each space group using CrystalFormer. On the other hand, we employ PyXtal [66] to generate crystal samples with the same stoichiometry in the same space groups. We then carry out structure relaxation using density functional (DFT) calculations.
| Space group | Crystal system | Converged structures | Retain symmetry | RMSD 1 () | 1 (eV/atom) | ||||
| CrystalFormer | PyXtal | CrystalFormer | PyXtal | CrystalFormer | PyXtal | CrystalFormer | PyXtal | ||
| (No. 2) | Triclinic | 46 | 67 | 45 | 67 | 1.929 | 2.038 | 1.034 | 0.913 |
| (No. 12) | Monoclinic | 55 | 72 | 53 | 67 | 1.846 | 2.235 | 1.233 | 1.660 |
| (No. 62) | Orthorhombic | 77 | 83 | 76 | 66 | 1.086 | 1.940 | 0.313 | 1.633 |
| (No. 139) | Tetragonal | 91 | 81 | 88 | 63 | 0.405 | 1.160 | 0.240 | 1.100 |
| (No. 166) | Trigonal | 83 | 74 | 80 | 71 | 1.028 | 2.183 | 0.352 | 2.489 |
| (No. 194) | Hexagonal | 97 | 77 | 96 | 60 | 0.355 | 2.129 | 0.324 | 4.100 |
| (No. 225) | Cubic | 98 | 96 | 95 | 92 | 0.076 | 0.518 | 0.214 | 0.483 |
| Average | 78.1 | 78.6 | 76.1 | 69.4 | 0.961 | 1.743 | 0.530 | 1.769 | |
1
Calculated on the converged structures.
Figure 7(a)-(g) shows the average energy difference to the energy of final structures versus DFT relaxation steps. We neglected the structures whose energy changes and energy change intervals per step during relaxations exceeded 10 eV/atom to eliminate the impact of erroneous steps. One sees that CrystalFormer samples generally reach lower energies in fewer relaxation steps. This is especially true for space groups with higher symmetries. The ability to initialize diverse and high-quality crystal structures enables one to discover more stable materials faster. Figure 7(h) shows the histogram of energy above the convex hull constructed by the Materials Project database. The dashed line denotes the criterion eV/atom [74] for selecting stable materials. Among these candidates, we found 34 and 12 relaxed structures with CrystalFormer and PyXtal initializations that are not contained in the MP-20 dataset. We summarize them in Table LABEL:table:samples and Table LABEL:table:pyxtal_samples of appendix C.
Table LABEL:table:700-stat lists detailed statistics of structure-relaxed samples in seven representative space groups. Overall, we find that the CrystalFormer generated structures are of higher quality, especially for those space groups with higher symmetry. This observation is supported by the fact that the DFT relaxation often retains the space group symmetry. The average root mean squared displacement (RMSD) computed for these converged structures is 0.961 and 1.743 for CrystalFormer and PyXtal initialization, respectively. In addition, the average energy above the convex hull also confirms the samples generated by CrystalFormer are indeed much closer to the DFT local minimum than PyXtal initialization.
IV.2 Structure-conditioned element substitution
Mutation of known crystals is a prominent approach to materials discovery. For example, one can employ a machine-learned force field to relax crystal structures [75, 5, 76, 77] after element substitutions. In the lens of generative modeling, the machine learning force field can be regarded as the energy-based model or Boltzmann machines. A potential drawback of exploring materials space with an energy-based model is the slow mixing or even ergodicity issue posed by the rough landscape of the potential energy surface. In this sense, element substitutions provide a variety of initial seeds, compensating for the limitation of energy-based exploration. Having an alternative measure of crystal likelihood other than the potential energy surface opens a way to employ the model likelihood as a guide for structure search.
Many crystal structures can be traced back to a few simple, highly symmetrical types. Numerous crystals share the same structural prototype but differ in composition, such as perovskite (ABX3), spinel (AB2X4), fluorite (AX2), and so on. Figure 8(a) shows double perovskite crystal structures A2BB’X6 which belong to the (No. 225) space group. There are hundreds of known double perovskites with significant interests in their semiconducting, ferroelectric, thermoelectric, and superconducting properties [78]. Finding more stable materials with this structure prototype using brute force enumeration and high-throughput calculation is a computationally demanding task [79]. We will generate new double perovskites with CrystalFormer and demonstrate its advantage of over standard element substitution methods.
Figure 8(a) shows the string representation of double perovskites. To generate candidates of double perovskites, we use CrystalFormer to carry out string in-filling tasks. Since the autoregressive sampling of the atoms is insufficient to take into account non-causal information in the sequence, we employ MCMC to sweep through the sequence and update chemical species and fractional coordinates [80]. The acceptance rate for these MCMC updates makes use of the marginalized probability for elements and fractional coordinates as the lattice parameter that appears at the end of the sequence can be integrated. Only after the MCMC sampling has been thermalized, we sample the lattice parameters autoregressive to account for the adjustment of the unit cell for given atoms and occupations. We use CrystalFormer to generate 100 candidates as the initial DFT relaxation.
As a comparison, we also employ the SubstitutionPredictorTransformation function [46] implemented in pymatgen [81] to perform element substitution for the crystals with double perovskite structures in the training dataset. The substitution probabilities come from data-mining of ICSD dataset [46]. After the substitution, we use DLSVolumePredictor [82] function of pymatgen to predict the volume of the structure. This lattice scaling scheme relies on data-mined bond lengths to predict the crystal volume of a given structure. To collect 100 candidates in the ionic substitution approach we have set the probability threshold of SubstitutionPredictorTransformation to 0.01, which is smaller than the typical values adopted in Ref. [48].
The RMSD computed for the DFT-relexed structures is 0.211 and 0.342 for CrystalFormer and ionic substitution [46], respectively. This observation confirms that the structures generated by CrystalFormer are closer to the DFT local minimum than ionic substitution [46]. Moreover, Figure 8(b) shows the histogram of energy above the convex hull of the Materials Project database. Overall, CrystalFormer and ionic substitution [46] found 9 and 3 double perovskites with eV/atom which are not contained in the MP-20 dataset, details in appendix D. The superior performance of CrystalFormer-guided MCMC is understandable since its likelihood takes into account the context of space group and atomic environment rather than marginal two-body correlation [46] in ionic substitution. The ionic substitution approach also shows two limitations in practical applications. First, some of the ions in the compound can not be substituted as they are missing in the probability table. Second, the approach relies on the calculability of the elements’ valence states.
As a final remark, although the discussion here focuses on generating crystals with given prototype structures, the generation of crystals with a given crystal lattice [83] is also feasible with CrystalFormer. This is because the crystal lattice can be straightforwardly expressed as constraints on the space group and occupied Wyckoff letters [57].
IV.3 Plug-and-play materials design
Finally, we demonstrate CrystalFormer’s ability to aid property-guided exploration of crystalline materials in a versatile and flexible manner. The trained CrystalFormer captures the space group conditioned crystal probability , which we treat as a prior probability for stable crystals. By combining it with a crystal property prediction model that provides the forward likelihood probability , one can carry out property-guided materials generation in a plug-and-play manner. According to Bayes’ rule, the posterior for crystals given property reads
By sampling from this posterior distribution, one can generate crystal samples with property guidance. Since the posterior probability Eq. (3) typically does not process autoregressive property with respect to , we carry out MCMC sampling to sample from the posterior distribution [84]. The plug-and-play feature makes designing crystalline materials in this way particularly appealing because it is possible to apply multiple conditions by simply adding log-likelihoods from multiple predictors. The framework applies to the inverse problem of solving cyrystal structures based on experimently observed diffraction spectra equally well [85, 86], where the goal is to simultaneously optimizing the matching probability to experimental observation and stability of the crystal.
Any property prediction model can used in conjunction with CrystalFormer for property-guided material generation. We train two crystal property prediction models using the MP-20 dataset and achieve mean absolute error (MAE) of 0.325 eV in band gap and MAE of 0.110 eV/atom in formation energy, which is comparable to the crystal graph convolutional neural network [87] trained on the same dataset [88]. We use the output of these two property prediction models as the forward probability of Eq. (3). More details are in appendix E.
Figure 9(a) demonstrates the controlled generation of materials with target band gap at and the formation of energy crystals [89, 90]. The conditional probability Eq. (3) contains both the model likelihood and the property regression MAE. Therefore, the generated samples will strike a balance between the two. To draw samples from the conditional probability distribution, we randomly generate a batch of 1000 crystal samples and sweep through the crystal sequence to update the atom species, fractional coordinates, and lattice parameters 222For simplicity of the Wyckoff sequence are kept unchanged in the MCMC sampling.. Monte Carlo update of chemical species so to achieve the desired properties can be regarded as a systematic data-informed way of carrying out cation-transmutation for materials inverse design [92]. After reaching equilibrium, the histogram of band gap and formation energy is centered around the target values, which is shifted significantly away from the value of unconditionally generated samples. On the other hand, the likelihood shown in Fig. 9(b) indicates these conditional-generated crystals are not typical samples with respect to the unconditional distributuion. Neverthless, they are still probable samples according to the crystal prior given by the CrystalFormer. Given a myriad of materials property prediction models developed over the years and the inconvenience of re-training or fine-tuning the foundational generative model [16], we envision the plug-and-play generation approach demonstrated here to be a scalable way for materials design. We have exposed an interface of CrystalFormer in the code repository [34] for users to plug in arbitrary conditioners for guided materials generation.
V Related works